When you think about the sums of their proper factors greater than 1, the numbers 48 and 75 have a special relationship. What is this relationship? show that the numbers 140 and 195 have this same special relationship.
Answer by stanbon(57410)

When you think about the sums of their proper factors greater than 1, the numbers 48 and 75 have a special relationship. What is this relationship? show that the numbers 140 and 195 have this same special relationship.
------------
48 = 2^2*3*4
sum of factors = 2+2+3+4 = 11
--------------
75 = 3*5^2
sum of factors = 3+5+5 = 13
==============
140 = 2^2*5*7
sum of factors = 2+2+5+7 = 16
--------------
195 = 3*5*13
sum of factors = 3+5+13 = 21
======================
If you see a pattern, good luck.
======================
Cheers,
Stan H.